# This work was supported by project from DOE-CESER

### v1

@@ 29 CTs Measurement and control Measurement High High In
S.No Equipment Funtionality Category Criticality during transients Vulnerability to extreme temperature Scope (in/out)
30 PTs Measurement and control Measurement High High In
31 Tap Changers Conversion Regulation Moderate High In
32 SVCs and VAR devices Voltage control Regulation Moderate High In
33 FACTs devices (STATCOM, SSSC, GUPFC, IPFC) Control and regulation Regulation High High In
34 Switchgear Panels Protection Protection High High In
35 Fuses Protection Protection High High In
36 HVDC converters Conversion and power transmission Operation High High In
37 Harmonic Filters Power quality Operation High High In
38 Disconnect switches and isolators Protection Control High High In
39 Capacitor Bank Control voltage level Regulation Moderate High In
40 batteries Energy storage Storage High High In
41 DLR System DLR System Measurement Moderate Minimal Out
@@ IV-A Insulators and Bushings
The primary function of insulators is to prevent flow of electric power away from the conductors. The key impact of failure of insulators results in faults where large currents can flow out of the electrical power grid leading to activation of protection systems (circuit breakers and fuses). The insulators are classified into several types depending on the location and function as shown in Fig. 8. However, the materials and their performance with respect or environmental conditions is expected to be similar [27, 28].
Most of the low voltage insulators typically in the distribution systems are some forms of polymers and ceramics like glass and porcelain [28]. Most overhead line insulators have different constituent materials like, rubber, insulating material, cement and iron couplings in their construction [29, 30, 31]. These joints between the various materials can get compromised under extreme heat. Literature reports extreme heat under conditions of wildfires, arc flash or flashovers. However, ambient conditions under conditions like heat domes and cycling of extreme temperatures can trigger failures of insulator structures [31]. There are other insulators used on busbars inside of switchgear and around auxiliary equipment that are more prone to failures due to extreme heat. These insulation options are usually among: Standoff insulators; epoxy power coating; heat-shrink tubes, and; insulating films [32]. The insulation over conductors under extreme heat coupled with high load currents tend to cause temperature rise to exceed beyond the safe operating ranges and tend to result in failures of insulators [33]. The thermal monitoring systems inside switchgear help with these high temperature related failures as described in subsection VI-K.
@@
Impact of Climate Change: While high temperatures can impact the mechanical components in the insulators that can lead to fatigue and failure, extreme cold conditions like polar vortex and winter storms can cause severe icing and snow depositions that can compromise the electrical function that can result in flash-overs and arcing across the insulators [34]. The extreme cold that occurs during polar vortexes results in icing and snow deposition on the outdoor insulators can lead to increased risk of leakage, arcing and flashovers that can cause temporary and permanent damage to the insulators [35].
Porcelain/ Ceramic insulators: Will have issues during extreme heat at the parts used for coupling to fixtures or conductors. These can have issues with degraded performance if resulting in a flashover or arcing due to snow or ice accumulation in extreme cold conditions during polar vortex and winter storms. Fig. 9 shows the snow deposition over insulators that has covered the entire surface of the insulator creating an easy path current leading to flashover [34, 35, 36].
Polymeric/Silicone Rubber (SR) Insulators: The polymeric insulators are made form a polymer that has better dielectric properties than a ceramic insulator and so are more compact than ceramic insulators. However, the polymeric insulators undergo permanent failure during a flashover unlike the ceramic insulators. This results in a higher risk due to extreme cold and ice storms where the insulators have ice deposition on the surface that can trigger flashovers and permanent damage.
@@
Cable Insulation: The cables and conductors in panels often have polymer based flexible insulation and these often tend to have failure modes in cases of extreme heat that usually occurs during heat domes. Figure 10 shows the insulation breakdown in cables and conductor insulation that usually results in failures resulting arc flash and equipment damage.
Figure 10: Failure and Monitoring for Cable Insulation from [38]
Existing Methods to Measure the Impacts: The key methods of monitoring the failures of insulators is mainly visual inspection. If the failures are large they tend to result is a large leakage current that can be detected by some relays depending on the sensing thresholds. There is discussion in the literature on power system insulator condition monitoring [39] that could prove to be useful for monitoring the condition and also for preventive maintenance.
Existing Methods to Mitigate the Impacts: Effective derating to consider the life and expected operating temperatures is a standard method to prevent failures during extreme heat. For managing extreme cold temperatures and severe icing, surface treatments like hydrophobic coatings that are typically used for switchgear panels and other power equipment could be explored for insulator and bushings [36]. While there are no specific methods discussed in the literature, additional planning and maintenance practices can be adopted to account for the changing climate to further consider derating the insulators that can ensure performance during the extreme climate conditions. An example of one such method is described in reference [40], that could prove to be effective in mitigating the impacts of insulator failures under extreme heat and cold or severe ice.
@@ IV-B Conductor
There are various types of power lines in power systems. Especially they are categorized as overhead and underground power lines. These two types of lines are present in both transmission and distribution power grids.
@@
Overhead Lines: Overhead lines are constructed by setting up conductors between poles. These poles can be built in any type of terrian and overhead lines are easier to repair damanges when compared to underground lines [41]. However, overhead lines are prone to extreme weather conditions due to their direct exposure to weather.
@@
Underground lines: Underground lines are constructed by laying down the conductors deep below the surface protected by conduits. Underground lines are less susceptible to weather but they are three to five times more costly when compared to overhead lines.
@@
Impact of Climate Change: The power lines are made of different conducting materials such as copper, aluminum, and steel. The conductors come in different configurations such as single core or multi core. Generally the material and structural configuration of conductor selection depends on the expected powerflow, losses, weather impact, etc. It is to be noted that the increase in temperature decreases the conductivity of a conductor irrespective of the material but reduction of conductivity varies for different materials. For example, primary material used for power lines are copper, aluminium, and steel. Copper is high conductive but expensive, while aluminium is $60$% of copper’s conductivity it is economical [42].
In early days, copper was used but it is replaced with aluminium or high pure aluminium wrapped around a steel core [43].
@@
However, due to global warming, there have been unusual/extreme changes in the temperatures than what has been historical observed, making it hard to ensure resiliency for power lines. In 2021, reports show that in Pacific Northwest, the temperatures were so high that the power lines were melted [44]. Furthermore, extreme cold has also stemmed disruptions in the winter [45]. Fig. 11 reports that the average annual total interruptions increase by 2 hours after excluding major events and it is reported that the primary reasons for major interruptions are due to weather followed by vegetation and maintenance issues [46]. Fig. 12 shows the U.S. states that have more frequent and more longer power interruptions, it is reported that U.S. has 18 weather related disasters that exceeded $1$ billion dollar damages and it includes winter storm Elliott that left Texas, Florida, Arkansas, Tennessee, West Virginia, and Wisconsin without power while also causing heavy snow fall of four feet in New York. [46]. It is also reported that the winter ice and snowstorms weigh down tree branches and power lines causing power outages. In 2023, reports indicate how extreme heat in combination with aging power grid infrastructure is more alarming situation for power grid reliability [47]. Furthermore, even underground cables can experience overheating when the loading in the power lines is underestimated but this can be mitigated forced cooling systems [48] albeit being an expensive solution.
@@
In extreme cold conditions, icing increases the weight of power lines. For example, trees can start to become so heavy that they may lean onto distribution lines, causing more outages, 1/2” of ice accumulation on power lines can add 500 pounds of extra weight [49]. Freezing rains in combination with steady winds also result in skipping rope like movement of power lines that causes outage. Furthermore, icing not only impacts overhead lines but it also impacts underground lines as ice follows the tree roots and builds up [49].
@@
Existing Methods to Measure the Impacts: The popular method to measure or monitor the impacts due to extreme weather on transmission lines is by using advanced weather monitoring tools [50]. Another solution is to directly monitor the transmission line temperature and its impact on dynamic line ratings [51]. In addition to conductor temperature and line ratings, this real-time measurement technology can also monitor the conductor clearance, conductor sag, and line current [52]. [53, 54] also shows how this temperature monitoring can be achieved based on Raman distributed optical fiber sensing technology.
Existing Methods to Mitigate the Impacts: There are several approaches to protecting the power lines in extreme temperature. One is more of planning solution where the risks due to climate change and historical climate risk are calculated, which are then used to identify the changes in power line ratings and other properties such as sag, clearance etc. [55, 56]. Another solution is to adapt underground cables more which are relatively less prone to extreme temperature when compared to overhead lines. However underground cables are also not fully immune to extreme temperature impacts as described in previous section. Furthermore, it is not practical to have very high voltage cables underground as they require additional cooling solutions making it very costly. Despite of these challenges, underground cables still help prevent outages due to extreme temperature at distribution level significantly. Other solution include rating the assets for different temperatures, continuous monitoring of heat, guarding against fires close to the power lines [57].
@@ IV-C Corona Ring
Corona rings are made of metal and used to distribute the electric field gradient and lower its value below the threshold to avoid corona discharge.
Impact of Climate Change: Temperature impacts corona rings in various ways. The corona onset voltage is the voltage at which corona discharge begins. [58, 59] presents the dependency of corona onset voltage and temperature i.e., as temperature increases, the corona onset voltage decreases. This is because rise in temperature increases the effective ionization coefficient. [60] discusses the impact of high temperatures on additional corona current due to factors such as gas molecule ionization, free electron detachment, and thermionic electron production. Increase in corona discharge has a number of negative impacts on the grid such as power loss, equipment damage, electromagnetic interference, insulation degradation, etc. [61] also shows that not only high temperatures impact (wear out faster) the corona rings but also low temperatures cause corona rings to operate below optimal efficiency. For this reason, corona rings are typically operated at ambient air or oil at room temperature (around $20\degree C$) even though they can be operated at wide range of temperatures.
Existing Methods to Measure the Impacts: The temperature of corona rings are primarily measured using infrared thermography to detect the heat emitted from the ring as a non-contact method [62].
Existing Methods to Mitigate the Impacts: Mitigation is a challenge due to outdoor nature and placement on the grid. The mitigation could come from design of corona rings to handle extreme environment conditions.
@@ V-A Fuses
Fuses are important protection components in distribution systems that carry electric power and are often classified as primary grid equipment. Reference [63] provides some fundamental information on parts and operation of fuses. The fuses operate to isolate faulty parts of the system very quickly when the fault currents are very huge. Fuses are designed such that when the large amount of fault current flows through it, the increase in the $I^{2}R$ losses cause excessive heating of the fuse element and in turn disconnects the circuit by blowing up. Therefore, it is clear that the fuse operation is heavily influenced by temperatures and extreme ambient temperatures can alter the normal operation of a fuse [64, 65, 66].
Typical life of a fuse is said to be anywhere between 20-30 years [67]. Over the lifespan of the fuse, there will be performance degradation. The key failure mode of the fuse is exposure to hard environmental conditions [68]. There are different types of fuses in the power systems as shown in the Fig. 13.
Figure 13: Extract: Types of Fuses in T&D Systems from [69]
Impact of Climate Change: Operating the fuse in extreme heat (similar to heat domes) or extreme cold conditions (polar vortex or icing) can lead to faster temperature cycling causing the expansion and contraction of the fuse element that is the most important part of the fuse. This repeated expansion and contraction can result is failure of the fuse [66, 70]. The other parts of the fuse that are sensitive to extreme temperatures are the contact elements and terminals. The parts inside the fuse also tend to degrade and are the other leading failure modes in fuses. Another common component in a fuse construction is the fuse enclosure and this is exposed to ambient operating conditions including heat domes, polar vortexes and icing. This enclosure has a limited function to just contain everything in a fuse internal to it but can transfer the ambient temperatures internal to the fuse. Extreme ambient conditions can impact the overall functioning as the enclosure cannot prevent ambient temperatures impacting the internal fuse temperatures. Often times heat domes and polar vortexes are short duration events (like a few consecutive days). During these short duration, derating the power system is important but, in the process, various controls and protection components like fuses and circuit breakers will need additional coordination of settings to prevent unwanted tripping leading to isolation of parts of the system that were not planned or intended. Fig. 14 below show a typical fuse on a distribution pole showing how it is exposed to the extreme operating conditions.
@@
The failure of a fuse mainly results in an open circuit and its impact on current carrying circuits is loss of power to downstream equipment/ loads. The failure of a fuse on a back-up circuit means the backup may not be functional causing serious reliability issues.
Existing Methods to Measure the Impacts: The impacts of a blown fuse are not damaging to the system components unlike other component failures, and therefore the key methods to identify if a Fuse is failed is usually visual inspection. There are some available products for the low voltage fuses and examples are shared in references [72] and [73]. A similar method could be used for medium and high voltage fuses in the systems.
Existing Methods to Mitigate the Impacts: During heat domes, the temperature rise and the heat in the ambient conditions affects the performance of the fuse and the fuse element. Therefore derating the fuse in these extremely hot conditions is a standard practice [70]. There is however a need to coordinate the fuse derating with the rest of the protection system. There is no literature that directly reports how the coordination is adopted in real world.

### v2

@@ 29 CTs measurement and control measurement high high yes
S.No Equipment Functionality Category Importance during grid events Vulnerability to extreme temperature Scoped-in (Yes/No)
30 PTs measurement and control measurement high high yes
31 tap changers conversion regulation moderate high yes
32 static VAR compensators and VAR devices voltage control regulation moderate high yes
33 flexible AC transmission devices (STATCOM, SSSC, GUPFC, IPFC) control and regulation regulation high high yes
34 switchgear panels protection protection high high yes
35 fuses protection protection high high yes
36 high-voltage direct current converters conversion and power transmission operation high high yes
37 harmonic filters power quality operation high high yes
38 disconnect switches and isolators protection control high high yes
39 dynamic line rating system calculate line capacity measurement moderate minimal no
@@ IV-A Insulators and Bushings
Insulators prevent the flow of electric power from conductors to supporting structures. When insulators fail, large currents can flow to ground, activating protection systems (circuit breakers and fuses). Insulators are classified into several types depending on location and function, as shown in Fig. 8. However, the materials and their performance with respect to environmental conditions is expected to be similar [27, 28].
Most low-voltage insulators (typically in distribution systems) are primarily made of polymers or ceramics like glass and porcelain [28], but these insulating materials are joined with other components. Overhead line insulators may include rubber, cement, iron couplings, and other materials [29, 30, 31]. Extreme heat can compromise the joints between materials. Literature reports extreme heat under conditions of wildfires, arc flash, or flashovers. Ambient conditions under heat domes and cycling of extreme temperatures can also trigger failures of insulator structures [31]. Insulators used on busbars inside of switchgear and around auxiliary equipment are more prone to failures due to extreme heat than insulators on overhead lines. These insulators include standoff insulators, epoxy powder coating, heat-shrink tubes, and insulating films [32]. Extreme heat coupled with high load currents can cause temperatures to rise beyond safe operating ranges, resulting in insulator failures [33]. The thermal monitoring systems inside switchgear help to prevent these high-temperature-related failures, as described in subsection VI-K.
@@
Impact of Extreme Temperature Operations: While high temperatures can affect the mechanical components of insulators, which can lead to fatigue and failure, extreme cold conditions like polar vortices and winter storms can cause severe icing and snow deposition that can compromise their electrical function [34]. Icing and snow deposition on outdoor insulators can lead to increased risk of leakage, arcing, and flashovers that can cause temporary and permanent damage to the insulators [35].
Porcelain/Ceramic Insulators: Extreme heat can cause issues with the parts used for coupling to fixtures or conductors. These insulators can also have issues with degraded performance, flashovers, or arcing due to snow or ice accumulation in extreme cold conditions during polar vortices and winter storms. In fig. 9, ice has covered the entire surface of the insulator and created an easy path for current, leading to flashover [34, 35, 36].
Polymeric/Silicone Rubber (SR) Insulators: Polymeric insulating materials have better dielectric properties than ceramics, so polymeric insulators are more compact. However, unlike ceramic insulators, polymeric insulators are highly vulnerable to damage from flashovers. This results in a higher risk from extreme cold and ice storms, because ice deposition that triggers a flashover can cause permanent failure of the insulator.
@@
Cable Insulation: Cables and conductors in panels often have polymer-based flexible insulation that can fail in extreme heat (e.g., during heat domes). Figure 10 shows how increased cable temperature can lead to insulation breakdown, in turn causing temperature to rise further. Cable insulation failure can result in arc flash and equipment damage.
Figure 10: Monitoring cable insulation temperature from [38]
Existing Methods to Measure the Impacts: Visual inspection is the main method of monitoring insulator failure. Large failures can result in a significant leakage current that can be detected by some relays, depending on the sensing thresholds. There is discussion in the literature on power system insulator condition monitoring [39] that could prove to be useful for for both monitoring and preventive maintenance.
Existing Methods to Mitigate the Impacts: Effective derating to consider the life and expected operating temperatures is a standard method to prevent failures during extreme heat. For managing extreme cold temperatures and severe icing, surface treatments like hydrophobic coatings that are typically used for switchgear panels and other power equipment could be explored for insulator and bushings [36]. While there are no specific methods discussed in the literature, additional planning and maintenance practices can be adopted to account for the changing climate, which could include further derating insulators to ensure performance during extreme climate conditions. Reference [40] describes one such method that could prove to be effective in predicting insulator failures under extreme heat and cold or severe ice.
@@ IV-B Conductor
There are various types of power lines in the T&D system, but they are essentially categorized as either overhead or underground power lines. These two types of lines are present in both transmission and distribution power grids.
@@
Overhead lines are constructed by setting up conductors between poles. These poles can be built in any type of terrain, and overhead lines are easier to repair than underground lines [41]. However, overhead lines are vulnerable to extreme weather conditions because of their direct exposure to weather.
@@
Underground lines are constructed by laying the conductors below the surface, protected by conduits. Underground lines are less susceptible to weather, but they are three to five times more costly than overhead lines [42].
@@
Power lines are made of different conducting materials such as copper and aluminum reinforced with steel cores. The conductors come in different configurations such as single core or multicore. The material and structural configuration depends on the expected powerflow, losses, weather impact, etc. Copper is highly conductive but expensive; while aluminum is $60$% of copper’s conductivity, it is much more economical [43]. Most power lines today are aluminum wrapped around a steel core [44]. Note that an increase in temperature decreases the conductivity of a conductor irrespective of the material, but reduction of conductivity varies for different materials. There are jumper cables used often to link ends of conductors that have similar properties as conductors but the mechanical joints for the jumper cables often need extra care to ensure their fitment and installation is accurate.
@@
Impact of Extreme Temperature Operations: Recent temperature extremes have made it more difficult to ensure resilience for power lines. In 2021, temperatures were so high in the Pacific Northwest that some power lines were reported to have melted [42]. Even underground cables can experience overheating when the loading in the power lines is underestimated, an effect which can be mitigated with forced cooling systems [45] but only at significant cost. Extreme cold has also caused disruptions in the winter [46]. Fig. 11 reports that the average annual total interruptions increase by 2 hours after excluding major events, and it is reported that the primary reasons for major interruptions are due to weather, followed by vegetation and maintenance issues [47]. Fig. 12 shows the U.S. states that have more frequent and more longer power interruptions. The U.S. has had 18 weather-related disasters that exceeded $1$ billion dollars in damages in 2022, including the winter storm Elliott, which left Texas, Florida, Arkansas, Tennessee, West Virginia, and Wisconsin without power and caused four feet of snowfall in New York [47].
@@
In extreme cold conditions, icing increases the weight of power lines and tree limbs. Trees can lean onto distribution lines, causing outages, and $1/2^{\prime\prime}$ of ice accumulation on power lines can add 500 pounds of extra weight [48]. Freezing rain in combination with steady winds also results in a skipping-rope-like movement of power lines (galloping) that causes outages. Icing not only affects overhead lines but also (much more rarely) underground lines, because ice follows tree roots and builds up [48].
@@
In addition these, extreme heal and cold operations make the jumper cables and their joints extremely vulnerable to become the weak links causing overall failure of the conductors. these can also become weak over the life of the conductor and the impact severity is increased on aged infrastructure operating under extreme temperatures.
Existing Methods to Measure the Impacts: Advanced weather monitoring tools are popular to measure or monitor the impacts of extreme weather on transmission lines [49]. Another solution is to directly monitor the transmission line temperature and its impact on dynamic line ratings [50]. In addition to conductor temperature and line ratings, this real-time measurement technology can also monitor the conductor clearance, conductor sag, and line current [51]. Refs. [52, 53] also show how this temperature monitoring can be achieved based on Raman distributed optical fiber sensing technology.
Existing Methods to Mitigate the Impacts: There are several approaches to protecting power lines against extreme temperatures. One is more informed planning; risks from climate change and historical climate risk are calculated and then used to identify the changes in power line ratings and other properties such as sag, clearance, etc. [54, 55]. Another solution is to adopt underground cables, which are less vulnerable to extreme temperatures than overhead lines. However, it is not practical to have very high voltage cables underground because they require additional cooling solutions, making them very costly. Despite of these challenges, underground cables still significantly help to prevent outages due to extreme temperature at the distribution level. Underground cables are also used at transmission level, but it is primarily done to navigate terrain and not common. Other solutions include rating the assets for different temperatures, continuous temperature monitoring, and guarding against fires close to the power lines [56].
@@ IV-C Corona Ring
Corona rings are made of metal and used in high-voltage transmission equipment to distribute the electric field gradient and lower its value below the corona threshold (i.e., the threshold above which electric current on a surface ionizes the surrounding air).
Impact of Extreme Temperature Operations: Temperature affects corona rings in various ways. The corona onset voltage is the voltage at which corona discharge begins. Ref. citezeng2023negative, huang2021research presents the dependency of corona onset voltage and temperature; as temperature increases, the corona onset voltage decreases. This is because rise in temperature increases the effective ionization coefficient. Ref. [59] discusses the impact of high temperatures on additional corona current due to factors such as gas molecule ionization, free electron detachment, and thermionic electron production. Increase in corona discharge has a number of negative impacts on the grid such as power loss, equipment damage, electromagnetic interference, insulation degradation, etc. Ref. [60] also shows that not only do high temperatures degrade corona rings, but low temperatures cause them to operate below optimal efficiency. For this reason, corona rings are typically operated at ambient air or oil at room temperature (around $20\degree$C) even though they can be operated at a wide range of temperatures.
Existing Methods to Measure the Impacts: The temperature of corona rings is primarily measured using infrared thermography to detect the heat emitted from the ring as a non-contact method [61].
Existing Methods to Mitigate the Impacts: Mitigation is a challenge due to outdoor nature and placement on the grid. The mitigation could come from design of corona rings to handle extreme weather conditions.
@@ V-A Fuses
Fuses are important protection components in distribution systems that carry electric power and are often classified as primary grid equipment. Reference [62] provides some fundamental information on parts and operation of fuses. The fuses operate to isolate faulty parts of the system very quickly when the fault currents are very large. Fuses are designed such that when the large amount of fault current flows through them, the increase in the $I^{2}R$ losses cause excessive heating of the fuse element and in turn disconnects the circuit by blowing up. Fuse operation is heavily influenced by temperatures and extreme ambient temperatures can alter the normal operation of a fuse [63, 64, 65].
Typical lifespan of a fuse is 20–30 years [66]; performance degradation occurs over that lifespan. The key failure mode of the fuse is exposure to hard environmental conditions [67]. There are different types of fuses in power systems, as shown in Fig. 13.
Figure 13: Types of fuses in T&D systems from [68]
Impact of Extreme Temperature Operations: Operating a fuse in extreme heat (as in heat domes) or extreme cold conditions (polar vortices or icing) can lead to faster temperature cycling, causing the expansion and contraction of the fuse element that is the most important part of the fuse. This repeated expansion and contraction can result is failure of the fuse [65, 69]. The other parts of the fuse that are sensitive to extreme temperatures are the contact elements and terminals. The parts inside the fuse also tend to degrade and are the other leading failure modes in fuses. Another common component in a fuse construction is the fuse enclosure, which is exposed to ambient operating conditions including heat domes, polar vortices, and icing. This enclosure has a limited function (simply to contain the fuse internals) but can transfer the ambient temperature to the fuse components. Extreme ambient conditions can impact the overall functioning because the enclosure cannot prevent ambient temperatures from affecting the internal fuse temperatures. Icing can cause mis-operation due to the malfunction of fuse cutout. Heat domes and polar vortices are often short-duration events (a few consecutive days). During these short durations, derating the power system is important, but in the process, various controls and protection components like fuses and circuit breakers will need additional coordination of settings to prevent unwanted tripping. Such tripping could lead to isolation of parts of the system that were not planned or intended. Fig. 14 shows a typical fuse on a distribution pole exposed to weather.
@@
The failure of a fuse mainly results in an open circuit and its impact on current-carrying circuits is loss of power to downstream equipment or loads. The failure of a fuse on a back-up circuit means the backup may not be functional, causing serious reliability issues.
Existing Methods to Measure the Impacts: Unlike other component failures, the impacts of a blown fuse are not damaging to the system components. Therefore, the key methods to identify if a fuse has failed is usually visual inspection. There are some available products for low-voltage fuses, and examples are shared in references [71] and [72]. A similar method could be used for medium- and high-voltage fuses.
Existing Methods to Mitigate the Impacts: During heat domes, the temperature rise and the heat in the ambient conditions affects the performance of the fuse and the fuse element. Therefore derating the fuse in these extremely hot conditions is a standard practice [69]. There is, however, a need to coordinate the fuse derating with the rest of the protection system. There is no literature that directly reports how the coordination is adopted in real world.